OpenStudy (anonymous):

Find the radius of a cylinder with a volume of 470.48 feet^3 and height of 6.5 feet.

jhonyy9 (jhonyy9):

do you can drawing it ?

jhonyy9 (jhonyy9):

hope you know sure formula for volume of a cylinder

jhonyy9 (jhonyy9):

V=area of base time height of cylinder area of base =pi*r^2 V=pi*r^2 *h 470,78 =3,14*r^2*6,5 so from above wrote you need calculi the r hope so much that you can calculi it easy good luck